1+cosx (2sinx-cosx)=sin^2x

Grade:11

2 Answers

Abhi99
9 years ago
what is rhs..? if it is (sinx)^2 then ans is x=n(pi)/2 ; n is an integer.
& if on rhs it is sin(2x) then ans is x=n(pi) ;n is an integer.
Jitender Singh
9 years ago
1 + cosx(2sinx – cosx) = sin2x
1 + 2sinx.cosx – cos2x = sin2x
1 + 2sinx.cosx = sin2x + cos2x
1 + 2sinx.cosx = 1
sin2x = 0
2x = n(pi)
x = n(pi) / 2